Held-to-Maturity Securities and Amortization of a Discount On January 1, 2019, Kelly Corporation acquired bonds with a face value of $500,000 for $483,841.79, a price that yields a 10% effective annual interest rate. The bonds carry a 9% stated rate of interest, pay interest semiannually on June 30 and December 31, are due December 31, 2022, and are being held to maturity.

Required: Prepare journal entries to record the purchase of the bonds and the first two interest receipts using the: 1. straight-line method of amortization 2. effective interest method of amortization
Answer #1

Part 1 ---- Journal Entries to record the purchase of the bonds and the first two interest receipts using Straight Line method of amortization

Semi Annual maturity period = 4*2 = 8
Issue Price of the Bonds = $483,841.79
Face Value / Maturity Value of the bonds =$500,000
Discount Received on Issue of the Bonds = $500,000 - $483,841.79 = $16,158.21
Amortization of Discount to be recognized on each payment = 16,158.21 / 8 = $2,019.78
Semi Annual Interest Receivable = Face Value x Stated Rate of Interest = $500,000 x 9% x ½ = $22,500
Date Account Titles and Explanation Debit Credit
Jan.1 Investment in Held to Maturity Debt Securities $483,841.79
Cash $483,841.79
June.30 Interest Receivable $22,500.00
Investment in Held to Maturity Debt Securities (Amortization Portion) $2,019.78
Interest Income $24,519.78
Dec.31 Interest Receivable $22,500.00
Investment in Held to Maturity Debt Securities (Amortization Portion) $2,019.78
Interest Income $24,519.78
Part2 ---- Journal Entries to record the purchase of the bonds and the first two interest receipts using Effective Interest Rate Method
Interest Revenue/Income = Carrying Value of the Investment in bonds at the beginning of period x Effective Interest Rate x Semi Annual ½
Interest Revenue on June 30.
Book/Carrying Value of Investment = $483,841.79
Interest Income = $483,841.79 x 10% x ½ = $24,192.09
Amortization of Discount = Interest Income – Cash Interest Receivable = $24,192.09 - $22,500 = $1,692.09
Interest Income on Dec.31
Book Value of Bond Investment = $483,841.79 + Amortization of Discount $1,692.09 = $485,533.88
Interest Income = $485,533.88 x 10% x ½ = $24,276.69
Amortization of Discount = Interest Income – Cash Interest Receivable = $24,276.69 - $22,500 = $1,776.69
Date Account Titles and Explanation Debit Credit
Jan.1 Investment in Held to Maturity Debt Securities $483,841.79
Cash $483,841.79
June.30 Interest Receivable $22,500.00
Investment in Held to Maturity Debt Securities (Amortization Portion) $1,692.09
Interest Income $24,192.09
Dec.31 Interest Receivable $22,500.00
Investment in Held to Maturity Debt Securities (Amortization Portion) $1,776.69
Interest Income $24,276.69